) Find P (0.45 < z < 2.1)

Answer by ewatrrr(24785)

 
Hi
Find P (0.45 < z < 2.1)
P (0.45 < z < 2.1)= P(z=2.1) - P(z =.45) = .9821 -.6736 = .3085
Note: Using Excel function: NORMSDIST(2.1)= .9821 and NORMSDIST(.45)= .6736